OpenStudy (anonymous):

solve for x 4x(x-5)-5x(x-4)=-1

OpenStudy (anonymous):

4x^2-20x-5x^2+20x=-1 -x^2=-1 x^2=1 \[x=\pm1\]

OpenStudy (anonymous):

so i have to put that symbol or can i leave it as 1

OpenStudy (anonymous):

Do you know what that symbol means?